OpenStudy (anonymous):

Problem for fun: Solve the system of equations: a + b + c = 4 ab + bc + ac = -17 abc = -60

OpenStudy (saifoo.khan):

substitution will work here.

OpenStudy (anonymous):

looks like the roots of a cubic :P

OpenStudy (anonymous):

P(x) = mx^3 +nx^2 +wx + z

OpenStudy (anonymous):

a = 5, b = 3, and c = -4

OpenStudy (anonymous):

actually, a, b, and c, all three could equal 5, 3, and -4.

OpenStudy (anonymous):

From the third equation,\[a \neq 0.\]Then, \begin{eqnarray*}a(b+c) + bc&=&-17 \\ a(4-a)+bc&=&-17 \\ a(4-a)-\frac{60}{a}&=&-17\\ a^3-4a^2-17a+60&=&0.\end{eqnarray*} Because this monic polynomial with integer coefficients, any rational solution (if it exists) must be an integer. This integer must also divide 60, so if a is rational, \[a \in \{\pm 1, \pm 2, \pm 3, \pm 4, \pm 5, \pm 6, \pm 10, \pm 15, \pm 20, \pm 30, \pm 60\}.\]Here I didn't find an elegant way, so I tested a few of these numbers and found a = -4 is a solution. Then, using long division (or Ruffini's algorithm or something similar) we can find the other two solutions, a = 3 and a = 5.

myininaya (myininaya):

very nice

OpenStudy (anonymous):

Nice indeed, thats what i was trying to get at. There was no difference in solving that system of solving the polynomial: \[x^{3} -4x^{2} -17x+60 = 0\] Thats because if i have a cubic polynomial with roots r, s, and t, then: \[(x - r)(x-s)(x-t) = x^{3}-(r+s+t)x^{2}+(rs+st+rt)x-(rst)\]